A 5.00 μF capacitor that is initially uncharged is connected in series with a 6.40 kΩ resistor and an emf source with E= 140 V negligible internal resistance.

Just after the circuit is completed, what is the voltage drop across the capacitor?

Just after the circuit is completed, what is the voltage drop across the resistor?

Just after the circuit is completed, what is the charge on the capacitor?

Just after the circuit is completed, what is the current through the resistor?
Answer #1

capacitance C=5 uF,

Resistance R=6.4*10^3 ohms

emf E=140V


a)

at t=0 sec,

voltage drop across capacitor, v=0


b)

voltage drop across resistor, vR=E=140 v


c)

charge on capacitor is zero


D)

current through the resistor i=VR/R

i=140/(6.4*10^3)

i=0.02187 A

i=21.9 mA